OpenStudy (anonymous):

Can someone help me with this? Integrate e^x/(e^x+4)

OpenStudy (lgbasallote):

hint!! let u = e^x do a u substitution ;)

OpenStudy (anonymous):

u = e^x + 4 du = e^x dx integrate 1/u du ... i think

OpenStudy (lgbasallote):

yup that's right @dedee :)

OpenStudy (lgbasallote):

do you know how to proceed now @Saltin8R ?

OpenStudy (anonymous):

not sure how this works do i answer or just help?

OpenStudy (anonymous):

Not exactly. Is there a certain rule for integrating 1/u? Coz that's just the same as u^-1, which would integrate to 1.

OpenStudy (lgbasallote):

you just help :) @dedee as much as possible ^_^

OpenStudy (anonymous):

yeah it should be listed in the back of your book or on a reference page

OpenStudy (lgbasallote):

yes there is a rule to 1/u it is the exception to the power rule \[\int \frac{1}{f(x)} = \ln |f(x)| + c\] got that @Saltin8R ?

OpenStudy (anonymous):

d/dx ln(x) = 1/x =)

OpenStudy (lgbasallote):

the power rule is \[\int f(x)^n = \frac{f(x)^{n+1}}{n+1}\] however..if n would be -1 then the denominator would be 0 garnering undefined..thus u^-1 is the exception

OpenStudy (anonymous):

yes, thanks guys :) I've got the correct answer :)

OpenStudy (lgbasallote):

congradumations ^_^